OpenStudy (anonymous):

A cheetah ran 300 feet in 2.92 seconds. What was the cheetah’s average speed in miles per hour?

OpenStudy (anonymous):

i know it's (300 ft/2.92 sec)(1 mi/ 5280 ft)(3600 sec/1 hour) = {300*1*3600 miles)/(2.92*5280*1 hour) = 91080000 miles/15417.6 hours = 70.05 mph

OpenStudy (anonymous):

but how would u rewrite it

OpenStudy (ghazi):

1m = 3.2ft 1 ft = 1/3.2 m 300 ft = \[\frac{ 300 }{ 3.2 }=93.75m\] now avg speed= 93.75/ 2.92 m/s